two aircraft P and Q are flying at the same speed 300 ms^-1 the direction along which p is flying is at right angle to the direction along which Q is flying
A)find the magnitude of the velocity of the aircraft P relative to aircraft QQ
B) find the direction of the velocity
1
Expert's answer
2019-12-04T08:41:26-0500

A)


"v=\sqrt{300^2+300^2}=424\frac{m}{s}"

B) The direction of the velocity is "45\degree" to the direction along which Q is flying and P is flying.